# Little Red Dots and Supermassive Black Hole Seed Formation in Ultralight Dark Matter Halos

### v1

@@ III Baryonic Collapse and Seed-Mass Scale inside ULDM Solitons
$T_{\rm core-cloud}\simeq\frac{3\mu m_{\rm H}}{16k_{\rm B}}v_{\rm in}^{2}\simeq\frac{3\mu m_{\rm H}}{8k_{\rm B}}\frac{GM_{b}(r_{c})}{r_{c}},$ (III.10)
where $v_{in}$ denotes the gas inflow velocity, $m_{H}$ is the hydrogen mass, and we take $\mu=1.22$ for definiteness. This leads to
$T_{\rm core-cloud}\simeq 96.33\ (1+z)^{\frac{3}{2}}\,\frac{\alpha^{2}}{m_{22}}\,{\rm K}.$ (III.11)
Radiation cooling down to the Jeans temperature of the core gas clouds is required to have a collapse. To estimate the Jeans temperature we proceed as follows. We assume that the gas clouds have an initial radius $r_{c}$ with the mass $M_{b}(r_{c})$ given in the above. We further assume that the clouds have a uniform density $\rho_{b}(r_{c})$ for $r<r_{c}$. The Jeans temperature can be obtained from the virial condition
$2E_{kinetic}+E_{potential}=0$ (III.12)
which becomes
$2\cdot\frac{3k_{B}T_{J}}{2}\frac{M_{b}(r_{c})}{\mu m_{H}}-\frac{3}{5}\frac{GM^{2}_{b}(r_{c})}{r_{c}}=0$ (III.13)
where $\mu m_{H}$ is the average mass of the particles. This leads to the Jeans temperature
$T_{J}=\frac{4\pi G\mu m_{H}\,r_{c}^{2}\rho_{b}(r_{c})}{15\,k_{B}}.$ (III.14)
With the above $r_{c}$ and $\rho_{b}(r_{c})$, one finds
$T_{J}\simeq 51.38\ (1+z)^{\frac{3}{2}}\,\frac{\alpha^{2}}{m_{22}}\,{\rm K}.$ (III.15)
For gravitational collapse to commence, the gas must first cool from $T_{\rm core\text{-}cloud}$ down to this temperature threshold. Once sufficient cooling has occurred, collapse can proceed. If $T_{J}<3\times 10^{4}{\rm K}$, the gas becomes prone to fragmentation, as discussed below. In contrast, if $T_{J}>3\times 10^{4}{\rm K}$, fragmentation is suppressed and direct collapse becomes possible. This requires
$\alpha^{2}\,\frac{(1+z)^{3/2}}{m_{22}\ }\ \gtrsim\ 583.9$ (III.16)
@@
$\alpha\ \gtrsim\ \alpha_{th}=24.16\frac{m^{1/2}_{22}}{(1+z)^{3/4}}$ (III.17)
@@
$M_{b}(r_{c})|_{\alpha=\alpha_{th}}=1.151\times 10^{6}m^{-1}_{22}\,{\rm M}_{\odot}.$ (III.18)
@@
This with the observational constraint for the minimum seed mass $10^{4}M_{\odot}\lesssim M_{bh}\lesssim 10^{5}M_{\odot}$ gives an interesting bound for the particle mass of ULDM
$1.151f_{bh}\times 10^{-21}eV\lesssim m\lesssim 1.151f_{bh}\times 10^{-20}eV,$ (III.19)
@@
which is consistent with the ULDM particle mass $m\simeq 10^{-22}eV$ inferred from galactic dynamics for $f_{bh}\simeq 0.1$. For $m\gg 10^{-22}eV$, SMBH seeds are too small in mass, whereas for $m\ll 10^{-22}eV$ their masses are excessively large and the seeds form too late (See Fig. 2). Gas clouds with masses below the threshold given in Eq. (III.18) fail to form black hole seeds and instead are expected to form star clusters. Interestingly, ULDM with the fiducial mass $m\simeq 10^{-22}eV$ seems to explain the minimum mass and the maximum mass of observed SMBHs. (See Fig. 3.)
From Eq. (III.17) and Eq. (III.3), one finds that the baryonic core size is smaller than $r_{c}|_{\alpha=\alpha_{th}}=4.884~{\rm pc}/m_{22}$, and the gas radius is smaller than $r_{vir}|_{\alpha=\alpha_{th}}\simeq 146.0~{\rm pc}/m_{22}$, which is consistent with LRD observations Guia et al. (2024). From Eq. (III.3) and Eq. (III.6), the mass inflow rate $\dot{M}=4\pi r_{c}^{2}\rho_{b}(r_{c})v_{in}$ is given by
$\dot{M}\simeq 2.310\times 10^{-3}\alpha^{3}m_{22}^{-3/2}(1+z)^{9/4}{\rm M}_{\odot}/{\rm yr}>32.58\,{\rm M}_{\odot}/{\rm yr},$ (III.20)
@@
where we used the fact $\alpha>\alpha_{th}$ for the last inequality. This strong inflow further ensures that the gas collapses monolithically to form a massive protostellar core, providing a dynamical pathway to SMBH seeds. Once SMBH seeds form, ULDM solitons can supply additional gravitational potential that enhances the Bondi accretion rate onto a growing black hole seed Chiu et al. (2025a), thereby enhancing the growth of SMBHs. Note that we do not consider angular momentum transfer and radiation feedback here.
@@
$M(z_{f})=M_{i}\exp\left[\frac{t(z_{f})-t(z_{i})}{t_{\rm Sal}}\right],$ (III.21)
@@
Figure 3: The mass spectrum of SMBHs as a function of redshift. The black horizontal line denotes the masses of SMBH seeds corresponding to $\alpha_{th}$. The blue dashed line and the purple dashed line denote the possible SMBH seed mass in Eq. (III.8) for $\alpha=1$ and $\alpha=\alpha_{max}$, respectively. However, objects located between the black horizontal line and the blue line cannot be genuine black hole seeds because the gas temperature is too low and fragmentation is unavoidable; instead, these systems are expected to form star clusters. According to Eq. (III.17) in our model, only gas clouds corresponding to the SMBH seeds in the gray region possess sufficient self-gravity and temperature to form a SMBH seed. The dashed horizontal red line denotes the maximum black hole mass $M_{bh}^{max}$ in Eq. (III.9) set by the stability of the soliton. Dots represent the observed masses of SMBHs from various observations Dewangan et al. (2008); Vestergaard and Osmer (2009); Willott et al. (2010); Mortlock et al. (2011); Larson et al. (2023); Natarajan et al. (2023). The red triangles denote XMM-Newton observations Dewangan et al. (2008), the green circles represent LBQS observations Vestergaard and Osmer (2009), the black diamonds denote CFHQS observations Willott et al. (2010). The red stars denote quasars with $z>7$ Mortlock et al. (2011); Larson et al. (2023); Natarajan et al. (2023), while the red star at $z=10.1$ is for the quasar in the galaxy UHZ1. The red disks denote the candidate SMBHs in LRDs Rusakov et al. (2026). We choose $m=10^{-22}eV$ and $f_{bh}=0.1$ here.
One can also determine the Jeans radius for baryon gas in terms of the average density $\rho_{b}$, and the temperature of the clouds as
$r_{J}=\left(\frac{15\,k_{B}T}{4\pi G\mu m_{H}\rho_{b}}\right)^{1/2}.$ (III.22)
Hence the baryon Jeans mass for a given density and temperature becomes
$M_{bJ}=\frac{4\pi}{3}\rho_{b}\,r_{J}^{3}=\left(\frac{375}{4\pi}\right)^{1/2}\left(\frac{k_{B}T}{G\mu m_{H}}\right)^{3/2}\rho^{-\frac{1}{2}}_{b}.$ (III.23)
During the collapse, we assume the adiabatic process. Then one has the temperature
$T=T_{i}\left(\frac{\rho_{b}}{\rho_{i}}\right)^{\gamma-1}$ (III.24)
as before where $T_{i}$ and $\rho_{i}$ denote the initial temperature and density of the collapse. As the collapse goes on, the density becomes larger and then temperature increases since $\gamma>1$ in general. Using this, one finds the updated baryon Jeans mass becomes
$M_{bJ}=M_{i}\left(\frac{\rho_{b}}{\rho_{i}}\right)^{\frac{3}{2}(\gamma-4/3)}.$ (III.25)
Therefore, if $\gamma<4/3$, $M_{bJ}<M_{i}=M_{b}(r_{c})$ and the fragmentation occurs as the collapse goes on. On the other hand, if $\gamma>4/3$, $M_{bJ}>M_{i}=M_{b}(r_{c})$ and the fragmentation is not allowed. When $T>3\times 10^{4}\,{\rm K}$, $\gamma\simeq 5/3$ and there is no fragmentation. If $T<3\times 10^{4}\,{\rm K}$, eventually the temperature passes through the range $T\sim 10^{4}$$3\times 10^{4}\,{\rm K}$ where $\gamma\simeq 1.1\sim 1.3$ Latif and Ferrara (2016). As a result, fragmentation into smaller stars occurs. When the collapse continues without fragmentation, a massive star with a mass on the order of $10^{5}M_{\odot}$ can form. This star can subsequently collapse into a black hole, as its lifetime is very short.
@@ IV Discussion
The deep gravitational potential of the soliton also leads to rapid gas inflow and shock heating up to $T_{\rm core\text{-}cloud}\gtrsim 3\times 10^{4}\,{\rm K}$, suppressing molecular hydrogen cooling and preventing fragmentation, consistent with the compact, hot, and ionized environments inferred for LRDs. Below the threshold mass, gas clouds generically fail to form black hole seeds and are instead expected to collapse into dense star clusters, possibly accounting for the observed coexistence of AGN-like and young stellar-dominated LRDs Carranza-Escudero et al. (2025), and potentially triggering the early formation of galaxies within a single physical framework. Furthermore, because the relevant Jeans scale and the soliton–halo relation favor the formation of dense cores at high redshift, our scenario predicts that SMBH seed formation is efficient at $z>5$, matching the redshift range where LRDs are observed by JWST. Finally, we demonstrate that standard Eddington-limited accretion is sufficient to grow these seeds to the observed masses by $z\sim 10$, without invoking super-Eddington accretion.
@@
The effects of angular momentum and radiative feedback are not included in the present semi-analytic treatment, although the solitonic potential is expected to promote rapid infall. While our model qualitatively reproduces key properties inferred for LRDs and SMBHs, including the characteristic mass scale, formation epoch, and the presence of hot, ionized gas, a more quantitative comparison with their number density and host-halo properties is still needed. Addressing these issues will require high-resolution simulations and improved observational constraints, and the present work should therefore be regarded as a starting point for more detailed studies linking the small-scale physics of ULDM to the observed population of high-redshift SMBHs.

### v2

@@ III Baryonic Collapse and Seed-Mass Scale inside ULDM Solitons
$T_{core}\simeq\frac{3\mu m_{\rm H}}{16k_{\rm B}}v_{\rm in}^{2}\simeq\frac{3\mu m_{\rm H}}{8k_{\rm B}}\frac{GM_{b}(r_{c})}{r_{c}},$ (III.10)
where $v_{in}\equiv\sqrt{2GM_{b}(r_{c})/r_{c}}$ denotes the gas inflow velocity, $m_{H}$ is the hydrogen mass, and we take $\mu=1.0$ for definiteness. This leads to
$T_{core}\simeq 78.96\ (1+z)^{\frac{3}{2}}\,\frac{\alpha^{2}}{m_{22}}\,{\rm K},$ (III.11)
which can easily exceed $10^{3}\,{\rm K}$ for $z>10$ and $\alpha\gg 1$.
From Eq. (III.6) one can obtain the number density of hydrogen atoms in the initial core
$n_{H}\simeq 0.2803\ \alpha^{4}(1+z)^{3}/{\rm cm}^{3}\,.$ (III.12)
These satisfy the relation
$T_{core}/n^{1/2}_{H}=q^{\prime}/m_{22}\,{\rm Kcm^{3/2}}$ (III.13)
with $q^{\prime}=78.96/0.2803^{1/2}\simeq 149.2$. For concreteness, in this paper we assume the direct collapse scenario in the reference Inayoshi and Omukai, 2012, where $H_{2}$ is collisionally dissociated and molecular cooling is suppressed. The gas remains on the efficient atomic cooling track at the temperature $T\simeq 8000\,{\rm K}$ and undergoes a nearly isothermal collapse, avoiding fragmentation and leading to the formation of a supermassive star, which can subsequently collapse into a black hole. From Eq. (9) and Eq. (10) of Ref. Inayoshi and Omukai, 2012, the the zone of no-return condition for direct collapse can be summarized as
$T_{core}\,n_{H}\geq q_{l}\,{\rm K/cm^{3}}\quad{\rm and}\quad T_{core}\,n_{H}^{0.1}\geq q_{r}\,{\rm K/cm^{0.3}},$ (III.14)
where $q_{l}=0.56\times 10^{8}$ and $q_{r}=0.5\times\sqrt{10}\times 10^{4}=1.581\times 10^{4}$. The two boundary lines meet at $T=(q_{r}^{10}/q_{l})^{1/9}{\rm K}=0.6377\times 10^{4}\,{\rm K}$ and $n_{H}=(q_{l}/q_{r})^{10/9}/cm^{3}=0.8782\times 10^{4}/cm^{3}$. If the initial values of $T_{core}$ and $n_{H}$ lie within the zone of no return, the direct collapse begins. Hence the direct collapse conditions are satisfied if
$T_{core}\geq(q_{l}(q^{\prime}/m_{22})^{2})^{1/3}{\rm K}=1.076\times 10^{4}\,{\rm K}/m^{2/3}_{22}\ {\rm and}$
$T_{core}\geq(q^{5}_{r}q^{\prime}/m_{22})^{1/6}{\rm K}\ =\ 0.7268\times 10^{4}\,{\rm K}/m^{1/6}_{22}.$ (III.15)
As both bounds are functions of $m_{22}$, only the larger of the two needs to be satisfied. The two lower bound temperatures become the same if $m_{22}=q^{\prime}q_{l}^{2/3}/q_{r}^{5/3}\equiv a_{0}=2.192$. Therefore, there are two cases.
If $m_{22}\leq a_{0}=2.192$, the collapse condition becomes $T_{core}\geq(q_{l}(q^{\prime}/m_{22})^{2})^{1/3}{\rm K}$. The threshold core temperature is of order $10^{4}\,{\rm K}$, which is high enough to dissociate $H_{2}$. In this case, from Eq. (III.11) direct collapse becomes possible if
$\alpha^{2}\,\frac{(1+z)^{3/2}}{m^{1/3}_{22}\ }\ \gtrsim\ (q_{l}{q^{\prime}}^{2})^{1/3}/78.96=136.3$ (III.16)
@@
$\alpha\ \gtrsim\ \alpha_{th}=(q_{l}{q^{\prime}}^{2})^{1/6}/\sqrt{78.96}\times\frac{m^{1/6}_{22}}{(1+z)^{3/4}}=11.67\times\frac{m^{1/6}_{22}}{(1+z)^{3/4}}$ (III.17)
@@
$M_{b}(r_{c})|_{\alpha=\alpha_{th}}=4.764\times 10^{4}\times(q_{l}{q^{\prime}}^{2})^{1/6}/\sqrt{78.96}\times m^{-4/3}_{22}\,{\rm M}_{\odot}=0.5562\times 10^{6}\times m^{-4/3}_{22}\,{\rm M}_{\odot}.$ (III.18)
Then, here and below, the minimum seed mass is given by
$M^{min}_{bh}\equiv f_{bh}M_{b}(r_{c})|_{\alpha=\alpha_{th}}.$ (III.19)
On the other hand, if $m_{22}\geq a_{0}$, the collapse condition becomes $T_{core}\geq(q^{5}_{r}q^{\prime}/m_{22})^{1/6}{\rm K}.$ In this case direct collapse becomes possible if
$\alpha^{2}\,\frac{(1+z)^{3/2}}{m^{5/6}_{22}\ }\ \gtrsim\ (q^{5}_{r}q^{\prime})^{1/6}/78.96=92.04$ (III.20)
or
$\alpha\ \gtrsim\ \alpha_{th}=(q^{5}_{r}q^{\prime})^{1/12}/\sqrt{78.96}\times\frac{m^{5/12}_{22}}{(1+z)^{3/4}}=9.594\times\frac{m^{5/12}_{22}}{(1+z)^{3/4}}$ (III.21)
which leads to
$M_{b}(r_{c})|_{\alpha=\alpha_{th}}=4.764\times 10^{4}\times(q^{5}_{r}q^{\prime})^{1/12}/\sqrt{78.96}\times m^{-13/12}_{22}\,{\rm M}_{\odot}=0.4571\times 10^{6}\times m^{-13/12}_{22}\,{\rm M}_{\odot}.$ (III.22)
@@
Combining these results with the observational constraint for the minimum seed mass $10^{4}M_{\odot}\lesssim M^{min}_{bh}\lesssim 10^{5}M_{\odot}$ yields interesting bounds for the particle mass of ULDM
$3.622~f_{bh}^{3/4}\lesssim m_{22}\lesssim 20.37~f_{bh}^{3/4},$ (III.23)
for the case 1 and
$4.067~f_{bh}^{12/13}\lesssim m_{22}\lesssim 34.07~f_{bh}^{12/13},$ (III.24)
@@
for the case 2, which are consistent with the ULDM particle mass $m\simeq O(10^{-22})eV$ inferred from galactic dynamics for $f_{bh}\simeq O(0.1)$. For $m\gg 10^{-22}eV$, SMBH seeds are too small in mass, whereas for $m\ll 10^{-22}eV$ their masses are excessively large and the seeds form too late (see Fig. 2). Interestingly, ULDM with the fiducial mass $m\simeq 10^{-22}eV$ seems to explain the minimum mass and the maximum mass of observed SMBHs. (See Fig. 3.) Cores of gas clouds with masses below the threshold given in Eq. (III.18) and in Eq. (III.22) fail to form black hole seeds and instead are expected to form star clusters.
Our model suggests the possible existence of LRDs. After SMBH seeds form inside the baryonic core, the remaining hot gas forms a cocoon surrounding the seeds. Let us consider the case 1 ($m_{22}\leq 2.192$) for example. From Eq. (III.17) and Eq. (III.3), one finds that the initial baryonic core size is smaller than
$r_{c}|_{\alpha=\alpha_{th}}=10.11~m^{-2/3}_{22}{\rm pc},$ (III.25)
and the radius of the whole gas cloud is smaller than
$r_{vir}|_{\alpha=\alpha_{th}}=c_{r}r_{c}|_{\alpha=\alpha_{th}}\simeq 303.3~m^{-2/3}_{22}{\rm pc},$ (III.26)
which is consistent with LRD observations Guia et al. (2024) implying an effective radius of gas clouds $\sim 100\,{\rm pc}$.
From Eq. (III.3) and Eq. (III.5), the mass inflow rate $\dot{M}=4\pi r_{c}^{2}\rho_{b}(r_{c})v_{in}=3\sqrt{2G}M_{b}(r_{c})^{3/2}/r_{c}^{3/2}$ is given by
$\dot{M}\simeq 2.308\times 10^{-3}\alpha^{3}m_{22}^{-3/2}(1+z)^{9/4}{\rm M}_{\odot}/{\rm yr}>3.669~m_{22}^{-1}{\rm M}_{\odot}/{\rm yr},$ (III.27)
@@
where we used the bound $\alpha>\alpha_{th}$ of the case 1 for the last inequality. This strong inflow ($\dot{M}\gg 0.1{\rm M_{\odot}/yr}$) further ensures that the gas collapses monolithically to form a massive protostellar core Umeda et al. (2016), providing a dynamical pathway to SMBH seeds. Once SMBH seeds form, ULDM solitons can supply additional gravitational potential that enhances the Bondi accretion rate onto a growing black hole seed Chiu et al. (2025a), thereby enhancing the growth of SMBHs. Note that we do not consider angular momentum transfer and radiation feedback here, for simplicity.
Figure 3: The mass spectrum of SMBHs as a function of redshift. The black horizontal line denotes the minimum masses of SMBH seeds $M^{min}_{bh}$ given in Eq. (III.19) corresponding to $\alpha_{th}$. We choose $m=10^{-22}eV$ and $f_{bh}=0.1$ here. The blue dashed line and the purple dashed line denote the possible SMBH seed mass in Eq. (III.8) for $\alpha=1$ and $\alpha=\alpha_{max}$, respectively. However, objects located between the black horizontal line and the blue line cannot be genuine black hole seeds because the gas temperature is too low and fragmentation is unavoidable; instead, these systems are expected to form star clusters. According to Eq. (III.18) in our model, only gas clouds corresponding to the SMBH seeds in the gray region possess sufficient self-gravity and temperature to form a SMBH seed. The dashed horizontal red line denotes the maximum black hole mass $M_{bh}^{max}$ in Eq. (III.9) set by the stability of the soliton. Dots represent the observed masses of SMBHs from various observations Dewangan et al. (2008); Vestergaard and Osmer (2009); Willott et al. (2010); Mortlock et al. (2011); Larson et al. (2023); Natarajan et al. (2023). The red triangles denote XMM-Newton observations Dewangan et al. (2008), the green circles represent LBQS observations Vestergaard and Osmer (2009), the black diamonds denote CFHQS observations Willott et al. (2010). The red stars denote quasars with $z>7$ Mortlock et al. (2011); Larson et al. (2023); Natarajan et al. (2023), while the red star at $z=10.1$ is for the quasar in the galaxy UHZ1. The red disks denote the candidate SMBHs in LRDs Rusakov et al. (2026).
@@
$M(z_{f})=M_{i}\exp\left[\frac{t(z_{f})-t(z_{i})}{t_{\rm Sal}}\right],$ (III.28)
@@ IV Discussion
The deep gravitational potential of the soliton also leads to rapid gas inflow and shock heating up to $T_{core}\sim 10^{4}\,{\rm K}$, suppressing molecular hydrogen cooling and preventing fragmentation, consistent with the compact, hot, and ionized environments inferred for LRDs. Below the threshold mass, gas clouds generically fail to form black hole seeds and are instead expected to collapse into dense star clusters, possibly accounting for the observed coexistence of AGN-like and young stellar-dominated LRDs Carranza-Escudero et al. (2025), and potentially triggering the early formation of galaxies within a single physical framework. Furthermore, because the relevant Jeans scale and the soliton–halo relation favor the formation of dense cores at high redshift, our scenario predicts that SMBH seed formation is efficient even at $z\geq 10$, explaining the early redshift range where LRDs are observed by JWST. Finally, we demonstrate that standard Eddington-limited accretion is sufficient to grow these seeds to the observed SMBH masses by $z\sim 10$, without invoking super-Eddington accretion.
@@
The effects of angular momentum and radiative feedback are not included in the present semi-analytic treatment, although the solitonic potential is expected to promote rapid infall. While our model qualitatively reproduces key properties inferred for LRDs and SMBHs, including the characteristic mass scale, formation epoch, and the presence of hot, ionized gas, a more quantitative comparison with their number density and host-halo properties is still needed. While the Press–Schechter approach provides a useful estimate, more accurate ULDM halo mass functions derived from simulations should be explored in future work. Addressing these issues will require high-resolution simulations Patrick et al. (2023) and improved observational constraints, and the present work should therefore be regarded as a starting point for more detailed studies linking the small-scale physics of ULDM to the observed population of high-redshift SMBHs.